Practice free →
Home › CBSE Class 11 (Commerce) › Mathematics › Probability › A coin is assigned P(H) = 0.3 under the axioms. …

A coin is assigned P(H) = 0.3 under the axioms. Then P(T) is

A0.3
B0.6
C0.7
D0.4
Answer & Solution
Correct answer: C. 0.7
With only two outcomes the probabilities sum to 1, so P(T) = 1 - 0.3 = 0.7.
Solve this in the app — CBSE Class 11 (Commerce) practice & 24k+ MCQs →
Related questions